OpenStudy (goformit100):

Check is my answer correct ? A student has to answer 10 questions, choosing atleast 4 from each of Parts A and B. If there are 6 questions in Part A and 7 in Part B, in how many ways can the student choose 10 questions?

OpenStudy (goformit100):

The possibilities are: 4 from Part A and 6 from Part B or 5 from Part A and 5 from Part B or 6 from Part A and 4 from Part B. Therefore, the required number of ways is 6C4×7C6+6C5×7C5+6C6×7C4= 105 + 126 + 35 = 266.

OpenStudy (goformit100):

@mikaela19900630

OpenStudy (anonymous):

Looks good to me!